What is the probability of picking any Ace from a full deck of 52 cards?
Because there are four aces in a 52-card deck, the probability is 4/52, or 1/13.
Animals probably evolved from?
Probably some sort of vaguely worm-like organism, sharing features of both animals and fungi (which is kinda our extremely distant relative). Before that, it's cells all the way down.
What is the probability of it raining in November 1st?
The probability is 1.
At any given time it is raining somewhere on the planet, so it is a certainty that it is raining on 1 November.
What is the probability that an offspring will be white?
The question depends on what the offspring is of!
What is the probability that 128 or fewer people have never played golf?
As there are millions of people who have no access to golf and millions more who are not physically capable of playing golf (through age, infirmity, parplegia, quadrapegia) etc. it is certain that the probability that 128 (or less) people have never played, or will ever play, golf - so a probability of 1.
If a random variable (RV) X is distributed Normally with mean m and standard deviation sthen
Z = (X - m)/s is the corresponding Normal variable which is distributed with mean 0 and variance 1. The distribution of X is difficult to compute but that for Z is readily available. It can be used to find the probabilities of the RV lying in different domains and thereby for testing hypotheses.
What is the probability of picking up number 15 from 1 to 20 numbers?
The probability, in a single random selection, is 1/20 or 0.05
If a dice is rolled 90 times what is the sum of the first and second roll expected to be on average?
If a DIE (not dice) is rolled 90 times, the expected value of the sum of the first and second rolls is 7 if you assume that the die is fair.
It does not matter how many times you roll the die, as long as it is at least 2.
What is a normal distribution graph called?
It has no special name - other than a normal (or Gaussian) distribution graph.
Is it true that you can improve your chances of having a boy or girl by doing certain things?
no.the chance to get a boy or a girl baby are 50%. The gender it is decided by speratozoa and it is casual.If spermatozoa it have x chromosome baby will be female, in case of Y chromosome spermatozoa baby will be a boy.
the only things that you can do it is to pay certain labs that will select spermtozoa by chromosome X or Y.in this case you can decide the sex of the child before you get pregnant.
50 times or above
What does nonreplacement mean in probability?
It refers to experiments where more than one tokens are randomly selected from a set of tokens (of different colours). If the the token is replaced after each selection, the probabilities remain constant whereas if the token is not replaced - as the question suggests - the probabilities change, depending on the outcome of the selection.
It refers to experiments where more than one tokens are randomly selected from a set of tokens (of different colours). If the the token is replaced after each selection, the probabilities remain constant whereas if the token is not replaced - as the question suggests - the probabilities change, depending on the outcome of the selection.
It refers to experiments where more than one tokens are randomly selected from a set of tokens (of different colours). If the the token is replaced after each selection, the probabilities remain constant whereas if the token is not replaced - as the question suggests - the probabilities change, depending on the outcome of the selection.
It refers to experiments where more than one tokens are randomly selected from a set of tokens (of different colours). If the the token is replaced after each selection, the probabilities remain constant whereas if the token is not replaced - as the question suggests - the probabilities change, depending on the outcome of the selection.
What is the probability of selecting march from all the months?
Total number of months we use to have = 12.
The month for required event is march. Quantity is one.
Therefore, required probability = 1/12.
There 52 cards in a deck What is the probability that you draw six black cards?
The probability that you draw any single black card would be:
(number of black cards)/(total number of cards)
and the probability to get 6 consecutive black cards would be:
(probability of drawing black) x (probability of getting another black)
and so on for 6 runs
Therefore (26/52) x (25/51) x (24/50) x (23/49) x (22/48) x (21/47) since there is one fewer black cards and one fewer total cards per draw.
(26/52) x (25/51) x (24/50) x (23/49) x (22/48) x (21/47) = .0113087788
= 1.13087788 %
What is the maximum number of foul ball possible in one at-bat?
There is no limit. The only real limit that would come even close is curfew. I remember as a kid (70's/80's), I believe it was the Tigers that the Tribe was playing, it was a really long extra inning game, and at either 1:00 or 2:00 am they stopped the game. (Both are Eastern time.) Though I believe, they finished out that game the next day before starting the new one. Kind of like an abbreviated double-header. So even that's not really a limit.
Don't know how rare, but as a 38 yr old woman, I unknowingly had mine for eight years and carried twins to full term before anything happened.
What are the chances of having twins after aborting twins?
It is the same for any other pregnancy. Abortion does not change the randomness of identical twins or the familiar tendency to have fraternal twins.
2300
The Central Limit Theorem (CLT) says no such thing! In fact, it states the exact opposite.
The CLT sets out the conditions under which you may use the normal distribution as an approximation to determine the probabilities of a variable X. If those conditions are not met then it is NOT OK to use the normal distribution.
It is 0.722... recurring.
What is the chance of rolling a two on one die?
1/6
The numerator is the number of outcomes you would find acceptable (in this case ONLY the number 2)
The denominator is the number of all possible outcomes (6)
25 marbles = 7 blue + 4 green + 8 red + 6 yellow
The ratio of blue to yellow marbles is 7:6
let
probability of killy goes to school=A= 0.39
probabilty of not going = A'
A'=1-A
P(A')=1-P(A)
P(A')=1-0.39
P(A')=0.61